40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 21 00:45:30 2015, Koji, Update, IOO, IMC fix PDRFsignal_cable.JPGQsignal_cable1.JPGQsignal_cable2.JPGQsignal_cable3.JPG
    Reply  Sat Nov 21 00:46:33 2015, Koji, Update, IOO, IMC OLTF IMC_OLTF.pdfPMC_noise_comparison.pdf
       Reply  Sun Nov 22 12:12:17 2015, Koji, Update, IOO, IMC OLTF 
    Reply  Sun Nov 22 07:09:01 2015, rana, Update, IOO, IMC fix 
       Reply  Sun Nov 22 12:07:09 2015, Koji, Update, IOO, IMC fix 
          Reply  Mon Nov 23 22:12:10 2015, Koji, Update, IOO, LO level check for the IMC demod board IMG_2137.JPGIMG_2138.JPG
             Reply  Wed Nov 25 04:24:21 2015, rana, Update, IOO, LO level check for the IMC demod board 
                Reply  Wed Nov 25 10:07:10 2015, Koji, Update, IOO, LO level check for the IMC demod board 
                   Reply  Wed Nov 25 16:46:10 2015, Koji, Update, IOO, LO level check for the IMC demod board D990511-40m_151123.pdf
                      Reply  Wed Nov 25 20:07:35 2015, rana, Update, IOO, LO level check for the IMC demod board 
                   Reply  Mon Nov 30 14:12:14 2015, ericq, Update, LSC, LO level check for the LSC RF distribution box 
                      Reply  Mon Nov 30 15:17:57 2015, Koji, Update, LSC, LO level check for the LSC RF distribution box 
Message ID: 11802     Entry time: Mon Nov 23 22:12:10 2015     In reply to: 11797     Reply to this: 11807
Author: Koji 
Type: Update 
Category: IOO 
Subject: LO level check for the IMC demod board 

In order to check the proper LO level, the IMC demod board was checked. As a short summary, -8dBm is the proper input for the IMC demod board. This was realized when the variable attenuator of the RF AM Stabilizer was set up be -7dB.


Initially, I tried to do the measurement using the extender board. But every board had the issue of +15V not working. After several extender boards were tried, I noticed that the current draw of the demod board burned the 15V line of the extender board.

Then I moved to the work bench. The signals were checked with the 10:1 probe. It's not properly the 50Ohm system, exactly to say.

I found that the LO signals at the mixers have huge distortion as it reaches the nominal 17dBm, and I wondered if ERA-5s were gone. Just in case I replaced the ERA-5s but didn't see any significant change. Then I thought it is due to the mixer itself. The mixer was removed and replaced with a 50Ohm SMD resister. Then the output of the last ERA-5 became sinusoidal, and the level was adjusted to be ~17dBm (4.52 Vpp) when the input power was measured to be -7.7dBm with the RF power meter. Once the mixer was reinstalled, it was confirmed that the waveform becase rectangular like, with the similar amplitude (4.42Vpp).

Now the module was returned to the rack. The RF level at the LO input was adjusted to be -8dBm by setting the attenuator level to be 7dBm.

Once the IMC is locked with this setting, the open loop transfer function was measured. The optical gain seemed almost unchanged compared with the recent nominal. The UGF and PM were measured to be 144kHz and 30deg.

Attachment 1: IMG_2137.JPG  7.169 MB  | Hide | Hide all
IMG_2137.JPG
Attachment 2: IMG_2138.JPG  7.253 MB  | Hide | Hide all
IMG_2138.JPG
ELOG V3.1.3-